OpenStudy (anonymous):

Write to numbers that have common factors of 1 3 and 5 explain how you found the numbers

OpenStudy (savvy):

factors of a number Xare those numbers which divide X completely leaving remainder = 0. if numbers a,b,c and d are given then, is abcd divisible by a,b,c and d seperately..???

OpenStudy (anonymous):

i dont get it

OpenStudy (savvy):

which part..???

OpenStudy (anonymous):

the whole thing...

OpenStudy (anonymous):

sorry i suck at math

OpenStudy (savvy):

its fine...tell me what is 8/2=??

OpenStudy (anonymous):

4

OpenStudy (savvy):

did 2 leave any remainder on dividing 8..???

OpenStudy (anonymous):

well 8/4=2 so no

OpenStudy (anonymous):

i mean 8/4=2...

OpenStudy (savvy):

so 2 is a factor of 8....

OpenStudy (anonymous):

ya

OpenStudy (savvy):

now as i said earlier... given 4 nos. a,b,c,d is abcd divisible by a..??? if yes then what is the quotient..???

OpenStudy (anonymous):

So you are looking for numbers with factors of fifteen. Those would be the multiples of fifteen, i.e., 15,30, 45, 60, 75,...